# Functionally Constrained Algorithm Solves Convex Simple Bilevel Problems

### v2

@@ Abstract
This paper studies simple bilevel problems, where a convex upper-level function is minimized over the optimal solutions of a convex lower-level problem. We first show the fundamental difficulty of simple bilevel problems, that the approximate optimal value of such problems is not obtainable by first-order zero-respecting algorithms. Then we follow recent works to pursue the weak approximate solutions. For this goal, we propose novel near-optimal methods for smooth and nonsmooth problems by reformulating them into functionally constrained problems.
@@ 1 Introduction
$\min_{\mathbf{x}\in\mathbb{R}^{n}}\;f(\mathbf{x})\quad\text{s.t.}\quad\mathbf{x}\in{\mathcal{X}}_{g}^{*}\triangleq\mathop{\arg\min}_{\mathbf{z}\in{\mathcal{Z}}}g(\mathbf{z}),$ (1)
@@
where $f^{*}$ is the optimal value of Problem (1) and $g^{*}$ is the optimal value of the lower-level problem. We highlight the asymmetry in $f$ and $g$ here. $f^{*}$ is the minimal in a constrained set, and hence it is possible that $f({\bf{x}})<f^{*}$ for some ${\bf{x}}\in{\mathbb{R}}^{n}$. On the other hand, $g^{*}$ is globally minimal and hence $g^{*}\leq g({\bf{x}})$ for any ${\bf{x}}\in{\mathbb{R}}^{n}$. Such asymmetry is natural as the role of $f,g$ are inherently asymmetrical for bilevel problems. We call such $\mathbf{\hat{x}}$ a $(\epsilon_{f},\epsilon_{g})$-absolute optimal solution.
@@
which we call a $(\epsilon_{f},\epsilon_{g})$-weak optimal solution. Much progress has been achieved towards this goal  [10, 24, 23, 18, 22, 17]. We note all the above algorithms fall in the class of zero-respecting algorithms (Assumption 3.4) and hence cannot obtain absolute optimal solutions.
@@
Then we further leverage the reformulation by Nesterov [19, Section 2.3.4] which relates the optimal value of Problem (4) to the minimal root of the following auxiliary function, where a discrete minimax problem defines the function value:
$\psi^{*}(t)=\min_{{\mathcal{Z}}}\left\{\psi(t,{\bf{x}})\triangleq\max\left\{f({\bf{x}})-t,\tilde{g}({\bf{x}})\right\}\right\}.$ (5)
@@
  • We propose a novel method FC-BiO that has near-optimal rates for finding $(\epsilon_{f},\epsilon_{g})$-weak optimal solutions of both nonsmooth and smooth problems.
@@ 2 Related work
In the literature, various methods [26, 25, 11, 3, 22, 17, 1, 13, 24, 23, 12, 6, 10, 18] have been proposed to achieve a $(\epsilon_{f},\epsilon_{g})$-weak optimal solution defined as Equation (3) . Below, we review the existing methods with non-asymptotic convergence. For ease of presentation, we state the results for $\epsilon_{f}=\epsilon_{g}=\epsilon$.
@@
Kaushik and Yousefian [13] proposed the averaging iteratively regularized gradient method (a-IRG) for convex optimization with variational inequality constraints, of which Problem (1) is a special case. a-IRG achieves the rate of ${\mathcal{O}}(1/\epsilon^{4})$. Shen et al. [24] proposed a method for solving Problem (1) with ${\mathcal{O}}(1/\epsilon^{3})$ complexity based on the online learning framework. When $f$ is Lipschitz and $g$ is smooth, Merchav and Sabach [18] proposed a gradient-based algorithm with ${\mathcal{O}}(1/\epsilon^{1/(1-\alpha)})$ complexity for any $\alpha\in(0.5,1)$. However, none of these methods can achieve the optimal rate of ${\mathcal{O}}(\epsilon^{-2})$.
@@
Samadi et al. [23] proposed the regularized accelerated proximal method (R-APM) with a complexity of ${\mathcal{O}}(1/\epsilon)$. Under the additional weak sharp minima condition on $g$, the complexity of R-APM improves to ${\mathcal{O}}\left(1/\sqrt{\epsilon}\right)$. However, this condition is often too strong and does not hold for many problems. Chen et al. [10] extended the result of [23] to the more general $\alpha$-Hölderian error bound condition. However, their method achieves the optimal rate only when $\alpha=1$, which reduces to the weak sharp minima condition. Jiang et al. [12] developed a conditional gradient type algorithm (CG-BiO) with a complexity of $\mathcal{{\mathcal{O}}}\left(1/\epsilon\right)$, which approximates $\mathcal{X}_{g}^{*}$ similar to the cutting plane approach. Later on, Cao et al. [6] proposed an accelerated algorithm with a similar cutting plane approach to achieve the rate of ${\mathcal{O}}(1/\sqrt{\epsilon})$ under the weak sharp minima condition. Recently, Wang et al. [27] reduced Problem (1) to finding the smallest $c$ such that the optimal value of the following parametric problem is $g^{*}$: $\min_{{\bf{x}}\in{\mathbb{R}}^{n}}g({\bf{x}}),~{}{\rm s.t.}~{}f({\bf{x}})\leq c$. They adopted a bisection method to find such a $c$. To solve this parametric problem, Accelerated Proximal Gradient method is applied on $g$ with projection operator onto the sublevel set ${\mathcal{F}}_{c}=\{{\bf{x}}~{}|~{}f({\bf{x}})\leq c\}$, which we call sublevel set oracles, which we call sublevel set oracles. This leads to an upper bound of $\tilde{\mathcal{O}}(1/\sqrt{\epsilon})$. Such an oracle is obtainable for norm-like functions such as $f({\bf{x}})=\frac{1}{2}\|{\bf{x}}\|^{2}$. However, projection onto sublevel sets may be hard for more general functions, such as MSE loss or logistic loss. Compared with prior works [27, 6, 10, 23], our proposed methods achieve the $\tilde{\mathcal{O}}(1/\sqrt{\epsilon})$ rate under standard assumptions, without relying on the weak sharp minima condition or the sublevel set oracles.
@@
Based on similar reformulation, Nesterov [19] has proposed algorithms for functionally constrained problems, of which Problem (4) is a special case: one for smooth problems in Section 2.3.5, and one for Lipschitz problems in Section 3.3.4. However, the first algorithm relies on the strong convexity of $f$ and $\tilde{g}$, which does not hold in our setups. Moreover, both algorithms require quadratic program (QP) sub-solvers, which are inefficient in large-scale problems: the first algorithm uses it to iteratively move $t$ to the root of $\psi^{*}(t)$, while the second algorithm uses it to project onto the sublevel set of a piecewise linear approximation of $\psi(t,{\bf{x}})$. In contrast, our proposed algorithms have the following advantages: (a) require no strong convexity assumption; (b) require no QP solvers that lead to additional computational cost; (c) have a unified framework for both smooth and Lipschitz problems.
@@ 3 Preliminaries
    The feasible set ${\mathcal{Z}}$ is convex and compact with radius $D=\sup_{{\bf{x}},{\bf{y}}\in{\mathcal{Z}}}\|{\bf{x}}-{\bf{y}}\|_{2}$.
The compactness assumption ensures that the subprocesses adopted in our method have a unified upper complexity bound (see Section 5.3). We note that other works involving bisection procedures, such as Wang et al. [27], may also need to address this issue to derive an explicit dependence on the distance term (although they did not state it formally in their paper). For unconstrained problems, if we know that the initial distance $\|{\bf{x}}^{*}-{\bf{x}}_{0}\|_{2}$ is upper bounded by $R$, we can simply take ${\mathcal{Z}}={\mathcal{B}}({\bf{x}}_{0},R)$.
@@
${\rm supp}({\bf{x}}_{t+1})\subseteq{\rm supp}({\bf{x}}_{0})\cup\left(\bigcup_{0\leq s\leq t}{\rm supp}\left(\partial f({\bf{x}}_{s})\right)\cup{\rm supp}(\partial g({\bf{x}}_{s}))\right).$ (6)
This assumption generalizes the standard definition of zero-respecting algorithm for single-level minimization problems [19, 7]. Many existing methods that incorporate a gradient step in the update for Problem (1) clearly fall within this class of algorithms, including those proposed by [10, 24, 23, 18, 22, 17], since the gradient step ensures that ${\bf{x}}_{k}\in{\bf{x}}_{0}+\mathrm{Span}\{\partial f({\bf{x}}_{0}),\partial g({\bf{x}}_{0}),\dots,\partial f({\bf{x}}_{k-1}),\partial g({\bf{x}}_{k-1})\}$. In the appendix, we show that the proposed algorithm in this paper and the conditional gradient type methods [12, 5] also satisfy the condition (6) when the domain is a Euclidean ball centered at ${\bf{x}}_{0}$ (see Proposition C.1 and Remark C.1).
The following concept of first-order zero-chain, introduced by Nesterov [19, Section 2.1.2], plays an essential role in proving lower bounds for zero-respecting algorithms. In our paper, we leverage the chain-like structure to show the intractability of finding absolute optimal solutions.
@@
${\rm supp}({\bf{x}}_{t+1})\subseteq\bigcup_{0\leq s\leq t}{\rm supp}\left(\nabla h({\bf{x}}_{s})\right),\;k\geq 1;\quad\mathbf{x}_{0}=\mathbf{0},$
@@
Definition 3.1 defines differentiable zero-chain functions. We can similarly define non-differentiable zero-chain, by requiring ${\rm supp}({\bf{x}}_{t+1})$ to be in $\bigcup_{0\leq s\leq t}{\rm supp}\left(\partial h({\bf{x}}_{s})\right)$, where $\partial h({\bf{x}})$ is a (possibly adversarial) subgradient of $h$ at ${\bf{x}}$.
@@ 5.2 Our proposed algorithms
We now present a unified framework applicable to both smooth and Lipschitz problems. The proposed algorithms nearly match the lower complexity bounds in both settings.
@@
We apply two steps of reformulation. First, we relax Problem (1) to Problem (4), where the constraint ${\bf{x}}\in{\mathcal{X}}_{g}^{*}$ is replaced by a relaxed functional constraint $\tilde{g}({\bf{x}})\triangleq g({\bf{x}})-\hat{g}^{*}\leq 0$. Denoting $\hat{f}^{*}$ as the optimal value of Problem (4), the following lemma holds:
@@
Such reformulation is introduced in Nesterov [19, Section 2.3] with the following characterization.
Lemma 5.2 (Nesterov [19, Lemma 2.3.4]).
@@
    $\psi^{*}(t)$ is continuous, decreasing, and Lipschitz with constant $1$.
@@
Algorithm 1 applies the bisection method within an initial interval $[\ell,u]$ which contains the smallest root, $\hat{f}^{*}$, for $N=\left\lceil\log_{2}\frac{u-\ell}{\epsilon/2}\right\rceil$ iterations. Similar to [27], the initial interval can be obtained by applying single-level first-order methods. The lower bound $\ell$ is obtained by solving the global minimum of the upper-level objective $f$, while $u=f(\hat{{\bf{x}}}_{g})$ serves as a valid upper bound, where $\hat{{\bf{x}}}_{g}$ is an approximate solution to the lower-level problem. Further details can be found in Appendix B.1. In each iteration, we set $t=\frac{\ell+u}{2}$. To approximate the function value of $\psi^{*}(t)$, we apply a first-order algorithm ${\mathcal{M}}$ to solve the discrete minimax problem (7). For the Lipschitz setting, we let ${\mathcal{M}}$ be the Subgradient Method (SGM, Algorithm 2) [4, Section 3.1]. For the smooth setting, we let ${\mathcal{M}}$ be the generalized accelerated gradient method (generalized AGM, Algorithm 3) [19, Algorithm 2.3.12]. These methods guarantee to find an approximate solution $\hat{\bf{x}}_{(t)}$ of Problem (7) such that
@@ 5.3 Subroutines and total complexity
Proposition 5.1 (Nesterov [19, Lemma 3.1.13]).
@@
The optimal first-order method for optimizing smooth objective functions is the celebrated Accelerated Gradient Method (AGM) [19, Section 2.2] proposed by Nesterov. In contrast to the Lipschitz setting, AGM cannot be applied to $\psi(t,{\bf{x}})=\max\{f({\bf{x}})-t,\tilde{g}({\bf{x}})\}$ directly when $f$ and $g$ are convex and smooth, as the smoothness condition no longer holds for $\psi(t,\cdot)$. However, Nesterov [19, Section 2.3] showed that by simply replacing the gradient step in standard AGM with the following gradient mapping (Nesterov [19, Definition 2.3.2])
@@
the optimal rate of $\mathcal{O}(\sqrt{L/\epsilon})$ can be achieved for Problem (7). Here $\{{\bf{x}}_{k}\},\{{\bf{y}}_{k}\}$ are the test point sequences and $L=\max\{L_{f},L_{g}\}$. Solving ${\bf{x}}_{k+1}$ for general discrete minimax problems (where the maximum is taken over potentially more than two objective functions, as studied in Nesterov [19, Section 2.2]), reduces to a quadratic programming (QP) problem and may not be efficiently solvable. However, we demonstrate that in our problem setup, ${\bf{x}}_{k+1}$ can be expressed in the form of a projection onto the feasible set ${\mathcal{Z}}$, or onto the intersection of ${\mathcal{Z}}$ and a hyperplane.
@@
Algorithm 3 Solve Problem (7) with Generalized AGM $({\bf{x}}_{0},D,t,K)$ [19, Algorithm 2.3.12]
@@
Lemma 5.5 (Nesterov [19, Theorem 2.3.5]).

### v3

@@ Abstract
This paper studies simple bilevel problems, where a convex upper-level function is minimized over the optimal solutions of a convex lower-level problem. We first show the fundamental difficulty of simple bilevel problems, that the approximate optimal value of such problems is not obtainable by first-order zero-respecting algorithms. Then we follow recent works to pursue the weak approximate solutions. For this goal, we propose a novel method by reformulating them into functionally constrained problems. Our method achieves near-optimal rates for both smooth and nonsmooth problems. To the best of our knowledge, this is the first near-optimal algorithm that works under standard assumptions of smoothness or Lipschitz continuity for the objective functions.
@@ 1 Introduction
$\min_{\mathbf{x}\in{\mathcal{Z}}}\;f(\mathbf{x})\quad\text{s.t.}\quad\mathbf{x}\in{\mathcal{X}}_{g}^{*}\triangleq\mathop{\arg\min}_{\mathbf{z}\in{\mathcal{Z}}}g(\mathbf{z}),$ (1)
@@
where $f^{*}$ is the optimal value of Problem (1) and $g^{*}$ is the optimal value of the lower-level problem ($\min_{z\in{\mathcal{Z}}}g(z)$). We highlight the asymmetry in $f$ and $g$ here. $f^{*}$ is the minimal in the constrained set ${\mathcal{X}}_{g}^{*}$, and hence it is possible that $f({\bf{x}})<f^{*}$ for some ${\bf{x}}\in{\mathcal{Z}}$. On the other hand, $g^{*}$ is globally minimal and hence $g^{*}\leq g({\bf{x}})$ for any ${\bf{x}}\in{\mathcal{Z}}$. Such asymmetry is natural as the role of $f,g$ are inherently asymmetrical for bilevel problems. We call such $\mathbf{\hat{x}}$ a $(\epsilon_{f},\epsilon_{g})$-absolute optimal solution.
@@
which we call a $(\epsilon_{f},\epsilon_{g})$-weak optimal solution. Much progress has been achieved towards this goal  [10, 25, 24, 19, 23, 18]. We note all the above algorithms fall in the class of zero-respecting algorithms (Assumption 3.4) and hence cannot obtain absolute optimal solutions, unless additional assumptions are made. (See Remark4.1 and Appendix D for further discussions.)
@@
where $\hat{g}^{*}$ is an approximate solution to the lower level problem $\min_{{\bf{x}}\in{\mathcal{Z}}}g({\bf{x}})$. Then we further leverage the reformulation by Nesterov [20, Section 2.3.4] which relates the optimal value of Problem (4) to the minimal root of the following auxiliary function, where a discrete minimax problem defines the function value:
@@
  • We propose a novel method FC-BiO that has near-optimal rates for finding $(\epsilon_{f},\epsilon_{g})$-weak optimal solutions of both nonsmooth and smooth problems. To the best of our knowledge, this is the first near-optimal algorithm that works under standard assumptions of smoothness or Lipschitz continuity for the objective functions. A comparison of previous results can be found in Section 2.
@@ 2 Related work
In the literature, various methods [27, 26, 12, 3, 23, 18, 1, 14, 25, 24, 13, 6, 10, 19] have been proposed to achieve a $(\epsilon_{f},\epsilon_{g})$-weak optimal solution to simple bilevel problems defined as Equation (3) . Below, we review the existing methods with non-asymptotic convergence. For ease of presentation, we state the results for $\epsilon_{f}=\epsilon_{g}=\epsilon$.
@@
Kaushik and Yousefian [14] proposed the averaging iteratively regularized gradient method (a-IRG) for convex optimization with variational inequality constraints, of which Problem (1) is a special case. a-IRG achieves the rate of ${\mathcal{O}}(1/\epsilon^{4})$. Shen et al. [25] proposed a method for solving Problem (1) with ${\mathcal{O}}(1/\epsilon^{3})$ complexity based on the online learning framework. When $f$ is Lipschitz continuous and $g$ is smooth, Merchav and Sabach [19] proposed a gradient-based algorithm with ${\mathcal{O}}(1/\epsilon^{1/(1-\alpha)})$ complexity for any $\alpha\in(0.5,1)$. However, none of these methods can achieve the optimal rate of ${\mathcal{O}}(\epsilon^{-2})$.
@@
Samadi et al. [24] proposed the regularized accelerated proximal method (R-APM) with a complexity of ${\mathcal{O}}(1/\epsilon)$. Under the additional weak sharp minima condition on $g$, the complexity of R-APM improves to ${\mathcal{O}}\left(1/\sqrt{\epsilon}\right)$. However, this condition is often too strong and does not hold for many problems. Chen et al. [10] extended the result of [24] to the more general $\alpha$-Hölderian error bound condition. However, their method achieves the optimal rate only when $\alpha=1$, which reduces to the weak sharp minima condition. Jiang et al. [13] developed a conditional gradient type algorithm (CG-BiO) with a complexity of $\mathcal{{\mathcal{O}}}\left(1/\epsilon\right)$, which approximates $\mathcal{X}_{g}^{*}$ similar to the cutting plane approach. Later on, Cao et al. [6] proposed an accelerated algorithm with a similar cutting plane approach to achieve the rate of ${\mathcal{O}}(\max\{1/\sqrt{\epsilon_{f}},1/\epsilon_{g}\})$, which can further be improved to ${\mathcal{O}}(1/\sqrt{\epsilon})$ under the additional weak sharp minima condition. Recently, Wang et al. [28] reduced Problem (1) to finding the smallest $c$ such that the optimal value of the following parametric problem is $g^{*}$: $\min_{{\bf{x}}\in{\mathbb{R}}^{n}}g({\bf{x}}),~{}{\rm s.t.}~{}f({\bf{x}})\leq c$. They adopted a bisection method to find such a $c$. To solve this parametric problem, Accelerated Proximal Gradient method is applied on $g$ with projection operator onto the sublevel set ${\mathcal{F}}_{c}=\{{\bf{x}}~{}|~{}f({\bf{x}})\leq c\}$, which we call sublevel set oracles. This leads to an upper bound of $\tilde{\mathcal{O}}(1/\sqrt{\epsilon})$. Such an oracle is obtainable for norm-like functions such as $f({\bf{x}})=\frac{1}{2}\|{\bf{x}}\|^{2}$. However, it may be computationally intractable for more general functions, such as MSE loss or logistic loss. It is a very strong oracle that is seldom used in the literature on optimization: the single-level optimization of a function $f$ using sublevel set oracles can be completed in ${\mathcal{O}}(\log(1/\epsilon))$ iterations of bisection procedure. Compared with previous work [28, 6, 10, 24], our proposed methods achieve the $\tilde{\mathcal{O}}(1/\sqrt{\epsilon})$ rate under standard assumptions, without assuming $f$ is a norm-like function or $g$ satisfies the weak sharp minima condition.
@@
Based on similar reformulation, Nesterov [20] has proposed algorithms for functionally constrained problems, of which Problem (4) is a special case: one for smooth problems in Section 2.3.5, and one for Lipschitz problems in Section 3.3.4. However, Nesterov’s algorithm for smooth problems relies on the strong convexity of $f$ and $g$, which does not hold in our bilevel setups. In this case, ${\mathcal{X}}_{g}^{*}$ would be a singleton, rendering the upper-level problem trivial. Our algorithm does not require the strong convexity assumption, and has a unified framework for both smooth and nonsmooth problems.
@@ 3 Preliminaries
    The feasible set ${\mathcal{Z}}$ is convex and compact with diameter $D=\sup_{{\bf{x}},{\bf{y}}\in{\mathcal{Z}}}\|{\bf{x}}-{\bf{y}}\|_{2}$.
The compactness assumption ensures that the subprocesses adopted in our method have a unified upper complexity bound (see Section 5.3). We note that other works involving bisection procedures, such as Wang et al. [28], may also need to address this issue to derive an explicit dependence on the distance term (although it is not stated formally in their paper). For unconstrained problems, if we know that the initial distance $\|{\bf{x}}^{*}-{\bf{x}}_{0}\|_{2}$ is upper bounded by $R$, we can simply take ${\mathcal{Z}}={\mathcal{B}}({\bf{x}}_{0},R)$.
@@
${\rm supp}({\bf{x}}_{k+1})\subseteq{\rm supp}({\bf{x}}_{0})\cup\left(\bigcup_{0\leq s\leq k}{\rm supp}\left(\partial f({\bf{x}}_{s})\right)\cup{\rm supp}(\partial g({\bf{x}}_{s}))\right).$ (6)
This assumption generalizes the standard definition of zero-respecting algorithm for single-level minimization problems [20, 7]. Many existing methods that incorporate a gradient step in the update for Problem (1) clearly fall within this class of algorithms, including those proposed by [10, 25, 24, 19, 23, 18], since the gradient step ensures that ${\bf{x}}_{k}\in{\bf{x}}_{0}+\mathrm{Span}\{\partial f({\bf{x}}_{0}),\partial g({\bf{x}}_{0}),\dots,\partial f({\bf{x}}_{k-1}),\partial g({\bf{x}}_{k-1})\}$. In the appendix, we show that the proposed algorithm in this paper and the conditional gradient type methods [13, 5] also satisfy the condition (6) when the domain is a Euclidean ball centered at ${\bf{x}}_{0}$ (see Proposition C.1 and Remark C.1), which suffices to establish the negative results, including the intractability results for absolute optimal solutions and lower complexity bounds for weak optimal solutions.
The following concept of first-order zero-chain, introduced by Nesterov [20, Section 2.1.2], plays an essential role in proving lower bounds for zero-respecting algorithms. In our paper, we leverage the chain-like structure to show the intractability of finding absolute optimal solutions.
@@
${\rm supp}({\bf{x}}_{k+1})\subseteq\bigcup_{0\leq s\leq k}{\rm supp}\left(\nabla h({\bf{x}}_{s})\right),\;k\geq 1;\quad\mathbf{x}_{0}=\mathbf{0},$
@@
Definition 3.1 defines differentiable zero-chain functions. We can similarly define non-differentiable zero-chain, by requiring ${\rm supp}({\bf{x}}_{k+1})$ to be in $\bigcup_{0\leq s\leq k}{\rm supp}\left(\partial h({\bf{x}}_{s})\right)$, where $\partial h({\bf{x}})$ is a (possibly adversarial) subgradient of $h$ at ${\bf{x}}$.
@@ 5.2 Our proposed algorithms
We now present a unified framework applicable to both smooth and Lipschitz problems. The proposed algorithms nearly match the lower complexity bounds in both settings up to logarithmic factors.
@@
We apply two steps of reformulation. First, we relax Problem (1) to Problem (4), where the constraint ${\bf{x}}\in{\mathcal{X}}_{g}^{*}$ is replaced by a relaxed functional constraint $\tilde{g}({\bf{x}})\triangleq g({\bf{x}})-\hat{g}^{*}\leq 0$ and $\hat{g}^{*}$ is an approximate solution to the lower level problem $\min_{{\bf{x}}\in{\mathcal{Z}}}g({\bf{x}})$. Denoting $\hat{f}^{*}$ as the optimal value of Problem (4), the following lemma holds:
@@
Such reformulation is introduced in Nesterov [20, Section 2.3] with the following characterization.
Lemma 5.2 (Nesterov [20, Lemma 2.3.4]).
@@
    $\psi^{*}(t)$ is continuous, decreasing, and Lipschitz continuous with constant $1$.
@@
Algorithm 1 applies the bisection method within an initial interval $[\ell,u]$ which contains the smallest root, $\hat{f}^{*}$, for $N=\left\lceil\log_{2}\frac{u-\ell}{\epsilon/2}\right\rceil$ iterations. Similar to [28], the initial interval can be obtained by applying single-level first-order methods. The lower bound $\ell$ is obtained by solving the global minimum of the upper-level objective $f$ over $x\in{\mathcal{Z}}$, while $u=f(\hat{{\bf{x}}}_{g})$ serves as a valid upper bound, where $\hat{{\bf{x}}}_{g}$ is an approximate solution to the lower-level problem. Further details can be found in Appendix B.1. In each iteration, we set $t=\frac{\ell+u}{2}$. To approximate the function value of $\psi^{*}(t)$, we apply a first-order algorithm ${\mathcal{M}}$ to solve the discrete minimax problem (7). For the Lipschitz setting, we let ${\mathcal{M}}$ be the Subgradient Method (SGM, Algorithm 2) [4, Section 3.1]. For the smooth setting, we let ${\mathcal{M}}$ be the generalized accelerated gradient method (generalized AGM, Algorithm 3) [20, Algorithm 2.3.12]. These methods guarantee to find an approximate solution $\hat{\bf{x}}_{(t)}$ of Problem (7) such that
@@ 5.3 Subroutines and total complexity
Proposition 5.1 (Nesterov [20, Lemma 3.1.13]).
@@
The optimal first-order method for optimizing smooth objective functions is the celebrated Accelerated Gradient Method (AGM) [20, Section 2.2] proposed by Nesterov. In contrast to the Lipschitz setting, AGM cannot be applied to $\psi(t,{\bf{x}})=\max\{f({\bf{x}})-t,\tilde{g}({\bf{x}})\}$ directly when $f$ and $g$ are convex and smooth, as the smoothness condition no longer holds for $\psi(t,\cdot)$. However, Nesterov [20, Section 2.3] showed that by simply replacing the gradient step in standard AGM with the following gradient mapping (Nesterov [20, Definition 2.3.2])
@@
the optimal rate of $\mathcal{O}(\sqrt{L/\epsilon})$ can be achieved for Problem (7). Here $\{{\bf{x}}_{k}\},\{{\bf{y}}_{k}\}$ are the test point sequences and $L=\max\{L_{f},L_{g}\}$. Solving ${\bf{x}}_{k+1}$ for general discrete minimax problems (where the maximum is taken over potentially more than two objective functions, as studied in Nesterov [20, Section 2.2]), reduces to a quadratic programming (QP) problem and may not be efficiently solvable. However, we demonstrate that in our problem setup, ${\bf{x}}_{k+1}$ can be expressed in the form of a projection onto the feasible set ${\mathcal{Z}}$, or onto the intersection of ${\mathcal{Z}}$ and a hyperplane. A similar subproblem also arises in Cao et al. [6, Remark 3.2]. When the structure of $\mathcal{H}$ is simple, such as when it is a Euclidean ball, the subproblem may admit a closed-form solution. Otherwise, Dykstra’s projection algorithm can be applied to solve it [11].
@@
Algorithm 3 Solve Problem (7) with Generalized AGM $({\bf{x}}_{0},D,t,K)$ [20, Algorithm 2.3.12]
@@
Lemma 5.5 (Nesterov [20, Theorem 2.3.5]).